A crate contains 62 oranges, 46 apples, and 94 pears. If 2 more of each type of fruit were added to the crate, each of the three types of fruit could be divided equally among a group of people. What is the greatest possible number of fruit in such group?
Answer by Theo(13342)

you have:
62 oranges
46 apples
94 pears.

2 more of each type are added to the crate.
you now have:
64 oranges
48 apples
96 pears

each of the fruits can be divided equally among a group of people.
what is the greatest possible number of fruit in such group?

i think you might have meant "what is the greatest number of people in such a group".

i think you are looking for the greatest common denominator.

it looks like the greatest common denominator is equal to 16.

since the greatest common denominator is 16, then i believe the greatest number of people in the group will be 16.

16 people will each get:

64/16 = 4 oranges each.
48/16 = 3 apples each.
96/16 = 6 pears each.